# Asymptotically good CSS-T codes exist

### v1

@@ Abstract
We give a new construction of binary quantum codes that enables the generation of a CSS-T code from any given CSS code. Using this construction, we prove the existence of asymptotically good binary CSS-T codes, resolving a previously open problem in the literature. Furthermore, we demonstrate that the same result holds for binary quantum low-density parity check CSS-T codes, and establish the existence of asymptotically good CSS codes that support any given $Z$ rotation transversally. Finally, we analyze the structure of the logical operators corresponding to certain non-Clifford gates supported by the quantum codes obtained from our construction.
Keywords— quantum code, CSS code, CSS-T code, asymptotically good codes, LDPC code, transversal gate
Introduction
Quantum error correction is a well-known and widely used technique for protecting quantum information from corruption by noise and, consequently, enabling the possibility of constructing fault-tolerant quantum computers. An $[\![n,k,d]\!]$ binary quantum error-correcting code encodes $k$ logical qubits of information into $n$ physical qubits and can correct up to $\lfloor\frac{d-1}{2}\rfloor$ errors. By itself, a quantum error correction code acts as a memory. However, operations over the logical level are required in order to obtain fault-tolerant quantum computers capable of realizing tasks that may exhibit quantum advantage. In this sense, logical operations are realized by manipulating the physical qubits encoding the information so that the desired logical operator is applied to the encoded logical qubits. Importantly, the operations applied to the physical qubits should preserve the code space, otherwise, the way in which the logical qubit is protected would essentially be lost. Furthermore, those logical operators must be induced in a fault-tolerant manner, implying that the overall quantum computation on the logical level is reliable enough.
@@ Introduction
In general, the Eastin–Knill theorem states that no universal gate set can be implemented on the logical level in a transversal manner for binary quantum error-correcting codes with $d>1$ [15]. Recently, a way to circumvent such a no-go theorem has been proposed under certain restrictions on the error model, but its application seems to be restricted to $d$-dimensional quantum systems [20]. There exist several universal quantum gate sets, with the Clifford+$T$ gate set being the most studied one [22]. Quantum circuits composed purely of Clifford gates are efficiently classically simulable as a result of the Gottesman–Knill theorem [18]. Implementing only transversal Clifford gates is both easier and less resource-intensive. In fact, finding codes accepting logical Clifford gates in a transversal manner is relatively straightforward. For example, any Calderbank–Shor–Steane (CSS) code constructed from a self-orthogonal binary cyclic code, or more generally doubly even self-orthogonal code, and its dual, admits transversal Clifford operators [12, 19, 36]. Nevertheless, those codes would lack the $T$-gate, implying that the quantum computations that can be realized would be classically simulable. A way to circumvent this problem is by means of magic state distillation and injection [7]. In this paradigm, magic state resources are injected into the “memory code” via a teleportation protocol resulting in the application of such operation on the code space in a fault-tolerant manner and the consumption of the magic resource [37]. The logical operation induced to the logical qubits will depend on the magic state resource injected, a common choice for the latter being |$\ket{T}$ magic states that result in the $T$ gate, which is the main non-Clifford gate discussed in this paper.
In order to apply the $T$ gate in a fault-tolerant manner, i.e., maintaining the error suppression level obtained by the code, magic states with a sufficient level of fidelity (of the order of the memory code) must be obtained. To do so, magic state distillation protocols such as the Bravyi–Haah protocols are used [4]. Such protocols make use of a certain amount of noisy magic states to distill a smaller number of magic states with higher fidelities. Magic state distillation protocols are closely connected to quantum error correction codes that admit the transversal implementation of a $T$-gate, i.e., such operation does not change the code space. These codes are commonly referred to as quantum codes that support transversal $T$.
Motivated by this application, numerous works in the literature have focused on constructing and characterizing stabilizer codes that support transversal $T$ gates; see, for instance, [1, 8, 9, 10, 32, 33]. In particular, [32] introduces CSS codes supporting transversal $T$ gates, known as CSS-T codes, which are the main object studied in this paper. The same authors demonstrate in [33] that CSS-T codes are optimal among all non-degenerate stabilizer codes supporting transversal $T$ gates.
However, a limitation of all currently known infinite families of CSS-T codes is that they exhibit either a vanishing rate or a vanishing relative distance. An intriguing open problem, proposed in [32], is to construct asymptotically good CSS-T codes, which simultaneously achieve a non-vanishing rate and relative distance.
@@
Motivated by the open problem mentioned above, we introduce a construction that produces a binary CSS-T code from any given binary CSS code. We employ this construction to prove the existence of asymptotically good binary CSS-T codes. Additionally, we show that our construction preserves the low-density parity check (LDPC) property, proving the existence of asymptotically good LDPC codes that are also CSS-T.
We also demonstrate that a generalization of our construction yields asymptotically good classes of binary codes that support transversal gates corresponding to any $Z$ rotation angle. Furthermore, we study the structure of the logical operators corresponding to the transversal $T$ gate, finer $Z$ rotations, and the $CCZ$ gate on our proposed CSS codes, which have the desired symmetry for these transversal gates.
@@
In Section 1, we set the notations and give definitions and known results on classical linear codes and quantum gates and codes. We also recall the definitions of CSS and CSS-T codes, which are the main focus of this paper. Section 2 presents our construction of CSS-T codes from CSS ones and their resulting parameters and properties. In particular, we prove here the existence of asymptotically good CSS-T codes. In Section 3, we discuss how to generalize our construction to support other transversal gates and examine their corresponding logical operators. Finally, in Section 4, we wrap up our work and present future research questions.
1 Preliminaries
In this section, we give a short introduction to classical and quantum error-correcting codes and their interplay. This allows us to fix the notations we will use for the rest of the paper. We refer the reader to [21] and to [17, 27, 35] for a detailed presentation of classical and quantum codes, respectively.
1.1 Classical linear codes
@@
The support of a vector $x\in\mathbb{F}_{2}^{n}$ is the set ${\textrm{S}upp}(x):=\{i\in\{1,\dots,n\}\,|\,x_{i}\neq 0\}$. The (Hamming) weight of a codeword $x\in C$ is defined by ${\textrm{w}t}(x):=|{\textrm{S}upp}(x)|$. If ${\textrm{w}t}(x)\equiv 0\pmod{2}$, then $x$ is called an even (weight) vector. A linear code $C\subseteq\mathbb{F}_{2}^{n}$ is called an even code if it only contains even vectors. For a set of codewords $S$ we use the notation ${\textrm{w}t}(S)\coloneqq\{{\textrm{w}t}(x)\mid x\in S,x\neq\mathbf{0}\}$, where $\boldsymbol{0}$ denotes the all-zeros vector. Finally, the minimum distance $d(C)$ of a code $C$ is defined as
$d(C)\coloneqq\min\{{\textrm{w}t}(x)\mid x\in C,\,x\neq\boldsymbol{0}\}.$
We will refer to a code of length $n$, dimension $k$ and minimum distance $d$ as an $[n,k,d]$ code.
The (Euclidean) dual of $C\subseteq\mathbb{F}_{2}^{n}$ is the linear code
$C^{\perp}\coloneqq\{y\in\mathbb{F}_{2}^{n}\mid x\cdot y=0\mbox{ for all }x\in C\},$
where $x\cdot y$ denotes the Euclidean inner product of the vectors $x$ and $y$. If $C$ has dimension $k$, then $C^{\perp}$ has dimension $n-k$. A code $C$ is said to be self-orthogonal if $C\subseteq C^{\perp}$, and self-dual if $C=C^{\perp}$.
The component-wise multiplication of two length $n$ vectors $x$ and $y$, also known as the Schur product, is defined as
@@
The Schur product of two codes $C,D\subseteq\mathbb{F}_{2}^{n}$ is then defined as
$C\star D\coloneqq{\textrm{S}pan}\{x\star y\mid x\in C,y\in D\}.$
The Schur product $C\star C$ is denoted $C^{\star 2}$. Note that for a binary linear code $C$, we have $x\star x=x$ for any $x\in C$, whence $C\subseteq C^{\star 2}$. We refer to [31] for an extensive survey on the Schur product of linear codes.
1.2 Quantum gates and quantum error correction
@@
where $v,u\in\mathcal{H}$ and $\overline{u_{i}}$ is the complex conjugate. A qubit of length one (or $1$-qubit) is an element of $\mathcal{H}$ of norm one. We show the set of one qubit Pauli operators by
$I=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$, $X=\begin{bmatrix}0&1\\ 1&0\end{bmatrix}$, $Y=\begin{bmatrix}0&-i\\ i&0\end{bmatrix}$, and $Z=\begin{bmatrix}1&0\\ 0&-1\end{bmatrix}$.
These operators form a basis for all $2\times 2$ matrices acting over $\mathcal{H}$. Hence an arbitrary one qubit error operator, i.e., $2\times 2$ unitary matrix over $\mathcal{H}$, can be represented as a linear combination of the mentioned matrices.
Let {|$\{\ket{0},\ket{1}\}$ be a basis for $\mathcal{H}$, where |$\ket{0}=\begin{bmatrix}1\\ 0\end{bmatrix}$ and |$\ket{1}=\begin{bmatrix}0\\ 1\end{bmatrix}$. An arbitrary state of a closed 1-qubit quantum system is defined by
|$\ket{\varphi}=\alpha\ket{0}+\beta\ket{1},$
[…]
Recall that the $Z$ rotation by the angle $\frac{2\pi}{2^{\ell}}$ for some $\ell\geq 0$ is given by the gate
@@
Then, the action of one qubit gate $\varGamma^{(\ell)}$ on the basis states is
$\begin{split}&\varGamma^{(\ell)}|0\rangle\longmapsto|0\rangle\ \text{and}\ \varGamma^{(\ell)}|1\rangle\longmapsto\mathrm{e}^{\frac{2\pi i}{2^{\ell}}}|1\rangle.\\ \end{split}$
If we fix $\ell=0,1,2,3$, then the $Z$ rotation $\varGamma^{(\ell)}$ is in correspondence to $I,Z,S$, and $T$, respectively.
For any $n\in\mathbb{N}$, we let $\mathcal{H}^{n}=\mathcal{H}\otimes\dots\otimes\mathcal{H}$ be the $2^{n}$-dimensional complex Hilbert space $\mathbb{C}^{2^{n}}$, where $\otimes$ denotes the usual tensor product of elements in $\mathbb{C}^{2}$. The inner product defined above for $\mathcal{H}$ generalizes to elements of $\mathcal{H}^{n}$. Qubits of length $n$ (or $n$-qubits) are elements of $\mathcal{H}^{n}$ of norm one. For the rest of this paper, we fix {|$\{\ket{a}\mid a\in\mathbb{F}_{2}^{n}\}$ to be a basis for $\mathcal{H}^{n}$.
1.3 From classical to quantum codes
We start by reviewing the construction of quantum CSS codes from a pair of classical codes. The following theorem summarizes the CSS construction of binary quantum codes, which is a commonly used technique in the literature.
@@ 1.3 From classical to quantum codes
Let $C_{2}\subseteq C_{1}$ be binary linear codes of length $n$ that define a binary quantum CSS code with parameters $[\![n,k,d]\!]$. Let $C_{1}=C_{2}\oplus{\textrm{S}pan}\{x_{1},x_{2},\ldots,x_{k}\}$, for some $x_{1},x_{2},\ldots,x_{k}\in C_{1}\setminus C_{2}$. Let $H$ be a $k\times n$ binary matrix having $x_{1},x_{2},\ldots,x_{k}$ as its rows. A common encoding approach for this binary quantum code is defined by ${\textrm{E}n}:\mathcal{H}^{k}\rightarrow\mathcal{H}^{n}$, where
${\textrm{E}n}\ket{u}=\frac{1}{\sqrt{|C_{2}|}}\sum_{x\in C_{2}}\ket{x+uH},$ (1)
where $u\in\mathbb{F}_{2}^{k}$. Note that the quantum state on the left-hand side of Equation (1), namely |$\ket{u}$, is a $k$-qubit state, while the right-hand side quantum states are $n$-qubit states. To distinguish them, from now on, we show the logical states using the notation |$\ket{u}_{L}$. Moreover, to simplify our computations, we discard the normalization factor. Each operation on such a quantum code is in the form of an $n$-qubit operation that preserves the code space, i.e., ${\textrm{E}n}(\mathcal{H}^{k})$.
In general, each error in an $n$-qubit operation can propagate within several of the code blocks. Therefore, a fault-tolerant approach is to investigate operations that can be decomposed into tensor product of $n$ single qubit gates, where each single error causes an error in one of the code blocks. Such $n$-qubit gates are called transversal gates. For instance, $T^{\otimes n}$ is an $n$-qubit transversal gate which will be called the transversal $T$ gate in this paper.
Recently, Rengaswamy et al. characterized all stabilizer codes in which transversal $T$ gate, and some other variants of it, preserve the code space [33]. In the literature, stabilizer codes with this property, specifically, those named triorthogonal, have been proposed for magic state distillation [4]. In [33], it also was shown that CSS codes are optimal among all non-degenerate stabilizer codes with this property. A restriction of their characterization to the CSS codes, commonly known as CSS-T codes, is stated below.
@@
A CSS code $(C_{1},C_{2})_{\mathrm{CSS}}$ is called a CSS-T code if the followings hold:

### v2

# Asymptotically good CSS-T codes and a new construction of triorthogonal codes

@@ Abstract
We propose a new systematic construction of CSS-T codes from any given CSS code using a map $\phi$. When $\phi$ is the identity map $I$, we retrieve the construction of [1] and use it to prove the existence of asymptotically good binary CSS-T codes, resolving a previously open problem in the literature, and of asymptotically good quantum LDPC CSS-T codes. We analyze the structure of the logical operators corresponding to certain non-Clifford gates supported by the quantum codes obtained from this construction ($\phi=I$), concluding that they always result in the logical identity. An immediate application of these codes in dealing with coherent noise is discussed. We then develop a new doubling transformation for obtaining triorthogonal codes, which generalizes the doubling construction presented in [2]. Our approach permits using self-orthogonal codes, instead of only doubly-even codes, as building blocks for triorthogonal codes. This broadens the range of codes available for magic state distillation.
Index Terms:
CSS code, CSS-T code, asymptotically good code, LDPC code, transversal gate, triorthogonal code, magic state distillation
I Introduction
Quantum error correction is a well-known and widely used technique for protecting quantum information from corruption by noise and, consequently, enabling the possibility of constructing fault-tolerant quantum computers. An $[\![n,k,d]\!]$ binary quantum error-correcting code encodes $k$ logical qubits of information into $n$ physical qubits and can correct up to $\lfloor\frac{d-1}{2}\rfloor$ errors. By itself, a quantum error correction code acts as a memory. However, operations over the logical level are required to obtain fault-tolerant quantum computers capable of realizing tasks that may exhibit quantum advantage. In this sense, logical operations are realized by manipulating the physical qubits encoding the information so that the desired logical operator is applied to the encoded logical qubits. Importantly, the operations applied to the physical qubits should preserve the code space; otherwise, how the logical qubit is protected would essentially be lost. Furthermore, those logical operators must be induced in a fault-tolerant manner, implying that the overall quantum computation on the logical level is reliable enough.
@@ I Introduction
In general, the Eastin–Knill theorem states that no universal gate set can be implemented on the logical level in a transversal manner for binary quantum error-correcting codes with $d>1$ [3]. Recently, a way to circumvent such a no-go theorem has been proposed under certain restrictions on the error model, but its application seems to be restricted to $d$-dimensional quantum systems [4]. There exist several universal quantum gate sets, with the Clifford+$T$ gate set being the most studied one [5]. Quantum circuits composed purely of Clifford gates are efficiently classically simulable as a result of the Gottesman–Knill theorem [6]. Implementing only transversal Clifford gates is both easier and less resource-intensive. In fact, finding codes accepting logical Clifford gates in a transversal manner is relatively straightforward. For example, any Calderbank–Shor–Steane (CSS) code constructed from a doubly-even code, and its dual, admits transversal Clifford operators [7, 8, 9]. Nevertheless, those codes would lack non-Clifford gates, implying that the quantum computations that can be realized would be classically simulable. A way to circumvent this problem is by means of magic state distillation and injection [10]. Magic state distillation protocols are closely connected to quantum error correction codes that admit the transversal implementation of a non-Clifford gate, i.e., such an operation does not change the code space. Note that for such a code to be useful for magic state distillation, the transversal implementation should not only preserve the codespace but also result in a non-Clifford operation at the logical level. A common family of such codes is that of triorthogonal codes [11].
Motivated by this application, numerous works in the literature have focused on constructing and characterizing stabilizer codes that support transversal $T$ gates; see, for instance, [12, 13, 14, 15, 16, 17]. In particular, [16] introduces CSS codes supporting physical transversal $T$ gates, known as CSS-T codes. The applicability of CSS-T codes to magic state distillation depends on the induced logical operators. Furthermore, quantum codes supporting physical transversal $T$ and finer rotations over the $Z$ axis that result in a trivial action on the logical space, i.e., logical identity, are useful to deal with coherent errors at the physical level [1]. A limitation of all currently known infinite families of CSS-T codes is that they exhibit either a vanishing rate or a vanishing relative distance [12, 2, 18]. An intriguing open problem, proposed in [16], is to construct asymptotically good CSS-T codes, which simultaneously achieve a non-vanishing rate and relative distance.
@@
Our main contribution is twofold. First, motivated by the open problem mentioned above, we introduce a construction that produces a binary CSS-T code from any given binary CSS code using a map $\phi$ respecting some properties. When $\phi$ is the identity map, we retrieve the construction of [1] and employ it to prove the existence of asymptotically good binary CSS-T codes. Additionally, we show that this construction preserves the low-density parity check (LDPC) property, proving the existence of asymptotically good LDPC codes that are also CSS-T. We also demonstrate that our method’s natural generalization yields asymptotically good CSS codes that support transversal gates corresponding to any $Z$ rotation angle. Furthermore, we study the structure of the logical operators corresponding to the transversal $T$ gate, finer $Z$ rotations, and the $CCZ$ gate on our proposed CSS codes, which have the desired symmetry for these transversal gates. In particular, we prove that all such transversal gates act as the logical identity operator when applied to the CSS codes from our construction. However, these codes have applications in mitigating coherent noise, and we discuss this application.
Secondly, we present a natural method for transforming certain CSS-T codes into triorthogonal codes. This result allows us to generalize the doubling construction discussed in [2] by permitting the use of general self-orthogonal codes, not necessarily self-dual, as input to the construction. This new perspective broadens the range of codes that can be used to generate triorthogonal codes, potentially enhancing the applicability of CSS codes in magic state distillation. In particular, we provide an example of a triorthogonal code that can be generated using our approach but cannot be obtained using the methods discussed in [2].
@@
In Section II, we set the notations and give definitions and known results on classical linear codes and quantum gates and codes. We also recall the definitions of CSS, CSS-T and triorthogonal codes, which are the main focus of this paper. Section III presents our construction of CSS-T codes from CSS ones and their properties. In particular, we prove here the existence of asymptotically good CSS-T codes. In Section IV, we discuss how to generalize our construction to support other transversal gates and examine their corresponding logical operators. Section V is devoted to our new construction of triorthogonal codes. Finally, in Section VI, we wrap up our work and present future research questions.
II Preliminaries
In this section, we give a short introduction to classical and quantum error-correcting codes and their interplay. This allows us to fix the notations used for the rest of the paper. We refer the reader to [19] and to [20, 21] for a detailed presentation of classical and quantum codes, respectively.
II-A Classical linear codes
@@
The support of a vector $x\in\mathbb{F}_{2}^{n}$ is the set ${\textrm{S}upp}(x):=\{i\in\{1,\dots,n\}\,|\,x_{i}\neq 0\}$. The (Hamming) weight of a codeword $x\in C$ is defined by ${\textrm{w}t}(x):=|{\textrm{S}upp}(x)|$. If ${\textrm{w}t}(x)\equiv 0\pmod{2}$, then $x$ is called an even (weight) vector. A linear code $C\subseteq\mathbb{F}_{2}^{n}$ is called an even code if it only contains even vectors. For a set of codewords $S$ we use the notation ${\textrm{w}t}(S)\coloneqq\{{\textrm{w}t}(x)\mid x\in S,x\neq\mathbf{0}\}$, where $\boldsymbol{0}$ denotes the all-zeros vector. Finally, the minimum distance $d(C)$ of a code $C$ is defined as $d(C)\coloneqq\min\{{\textrm{w}t}(x)\mid x\in C,\,x\neq\boldsymbol{0}\}.$ We will refer to a code of length $n$, dimension $k$ and minimum distance $d$ as an $[n,k,d]$ code.
The (Euclidean) dual of $C\subseteq\mathbb{F}_{2}^{n}$ is the linear code $C^{\perp}\coloneqq\{y\in\mathbb{F}_{2}^{n}\mid x\cdot y=0\mbox{ for all }x\in C\},$ where $x\cdot y$ denotes the Euclidean inner product of the vectors $x$ and $y$. If $C=C^{\perp}$, the code is called self-dual. The component-wise multiplication of $x,y\in\mathbb{F}_{2}^{n}$, or the Schur product, is
@@
The Schur product of two codes $C,D\subseteq\mathbb{F}_{2}^{n}$ is then defined as $C\star D\coloneqq{\textrm{S}pan}\{x\star y\mid x\in C,y\in D\}.$ The Schur product $C\star C$ is denoted $C^{\star 2}$. One can easily see that $C\subseteq C^{\star 2}$. We refer to [22] for an extensive survey on the Schur product of codes.
II-B Quantum gates and quantum error correction
@@
where $v,u\in\mathcal{H}$ and $\overline{u_{i}}$ is the complex conjugate. A qubit of length one (or $1$-qubit) is an element of $\mathcal{H}$ of norm one. We show the set of one-qubit Pauli operators by $I,X,Y,$ and $Z$. Let {|$\{\ket{0},\ket{1}\}$ be a basis for $\mathcal{H}$, where |$\ket{0}=\begin{bmatrix}1\\ 0\end{bmatrix}$ and |$\ket{1}=\begin{bmatrix}0\\ 1\end{bmatrix}$. An arbitrary state of a closed 1-qubit quantum system is defined by |$\ket{\varphi}=\alpha\ket{0}+\beta\ket{1},$ where $\alpha,\beta\in\mathbb{C}$ and $|\alpha|^{2}+|\beta|^{2}=1$. The action of each of the Pauli operators $P$ on |$\ket{\varphi}$ is defined by $P\ket{\varphi}$. Recall that the $Z$ rotation by the angle $\frac{2\pi}{2^{\ell}}$ for some $\ell\geq 0$ is given by the gate
@@
Then, the action of $\varGamma^{(\ell)}$ on the basis states is $\varGamma^{(\ell)}|0\rangle\longmapsto|0\rangle$ and $\varGamma^{(\ell)}|1\rangle\longmapsto\mathrm{e}^{\frac{2\pi i}{2^{\ell}}}|1\rangle.$ If we fix $\ell=0,1,2,3$, then the $Z$ rotation $\varGamma^{(\ell)}$ is in correspondence to $I,Z,S$, and $T$, respectively.
For any $n\in\mathbb{N}$, we let $\mathcal{H}^{n}=\mathcal{H}\otimes\dots\otimes\mathcal{H}$ to be the $2^{n}$-dimensional complex Hilbert space $\mathbb{C}^{2^{n}}$, where $\otimes$ denotes the usual tensor product of elements in $\mathbb{C}^{2}$. Qubits of length $n$ (or $n$-qubits) are elements of $\mathcal{H}^{n}$ of norm one. For the rest of this paper, we fix {|$\{\ket{a}\mid a\in\mathbb{F}_{2}^{n}\}$ to be a basis for $\mathcal{H}^{n}$.
II-C From classical to quantum codes
We start by reviewing the construction of binary quantum CSS codes from a pair of classical codes.
@@ II-C From classical to quantum codes
Let $C_{2}\subseteq C_{1}$ be binary linear codes of length $n$ that define a binary quantum CSS code with parameters $[\![n,k,d]\!]$. Let $C_{1}=C_{2}\oplus{\textrm{S}pan}\{x_{1},x_{2},\ldots,x_{k}\}$, for some $x_{1},x_{2},\ldots,x_{k}\in C_{1}\setminus C_{2}$. Let $H$ be a $k\times n$ binary matrix having $x_{1},x_{2},\ldots,x_{k}$ as its rows. A common encoding approach for this binary quantum code is defined by ${\textrm{E}n}:\mathcal{H}^{k}\rightarrow\mathcal{H}^{n}$, where for $u\in\mathbb{F}_{2}^{k}$ we have
${\textrm{E}n}\ket{u}=\frac{1}{\sqrt{|C_{2}|}}\sum_{x\in C_{2}}\ket{x+uH}.$ (1)
Note that the quantum state |$\ket{u}$ on the left-hand side of Equation (1) is a $k$-qubit state, while the right-hand side quantum states are $n$-qubit states. To distinguish them, from now on, we show the logical states using the notation |$\ket{u}_{L}$. Moreover, to simplify our computations, we discard the normalization factor. Each operation on such a quantum code is in the form of an $n$-qubit operation that preserves the code space, i.e., ${\textrm{E}n}(\mathcal{H}^{k})$.
Recently, Rengaswamy et al. characterized all stabilizer codes in which the transversal $T$ gate, and some other variants of it, preserve the code space [17]. In [17], it was also shown that CSS codes are optimal among all non-degenerate stabilizer codes with this property. A restriction of the characterization from [17] to CSS codes, commonly known as CSS-T codes, is stated below.
@@
A CSS code $(C_{1},C_{2})_{\mathrm{CSS}}$ is called a CSS-T code if the following hold: